OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use sqrt = 3.14. d = 12 A = sq. units i got 18.84 but its rong

OpenStudy (stormfire1):

3.14*6^2 = 113.04 square units

OpenStudy (hoblos):

A = 3.14 * 6^2 =113.04

OpenStudy (stormfire1):

You have to SQUARE the radius :)

OpenStudy (anonymous):

ok so what if we change it up and the d=10

OpenStudy (anonymous):

????

OpenStudy (stormfire1):

3.14*5^2=?

OpenStudy (stormfire1):

Since 5 is 1/2 of 10...

OpenStudy (anonymous):

is it 31.4

OpenStudy (anonymous):

???

OpenStudy (stormfire1):

no...it's 3.14 * (5^2) = 78.5

OpenStudy (stormfire1):

You can't use the diameter...you have to use the radius in the equation. Again, radius = 1/2 of the diameter.

OpenStudy (anonymous):

ok thnx